OpenStudy (anonymous):

Find the equation of the line with slope m=-1 passing through the point (3,-2) ?!

OpenStudy (radar):

slope m=-1 given. you also know \[m =\Delta y/\Delta x\] then m=(y -(-2))/(x-3) -1=(y+2)/(x-3) y=-x+1